OpenStudy (anonymous):

What is the 24th term of the arithmetic sequence 10, 16, 22, … ?

OpenStudy (abb0t):

148

OpenStudy (abb0t):

a(n) = a(1) + (n - 1)d n = which term d = common difference a(1) = first term a(n) = the term you are looking for

OpenStudy (abb0t):

a(24) = 10 +(24 - 1)(6)

OpenStudy (anonymous):

thnx, bro